The valence electrons in nitrogen would experience a------effective nuclear charge. Be has four protons, and N has seven protons. Both atoms have ---- core electrons that predominately contribute to the shielding, while the valence electrons will contribute a slight shielding effect. So Be has an effective nuclear charge of slightly more than -----, and N has an effective nuclear charge of slightly more than 5+ .

greater

2+

3+

lesser

two

1+

four

three
Answer #1

Be:: Z= 4, 1s2 2s2
effective nuclear charge can be calculated using following relation
Zeff = Z- S where S= core electrons = 2
= 4-2 = 2
So Zeff will be slightly more than 2+
************

N ::Z=7, 1s2 2s2 2p3
Zeff = Z- S where S= core electrons = 2
Zeff = 7-2 = 5
So Zeff will be slightly more than 5+
********************
SO answer is
The valence electrons in nitrogen would experience a GREATER
effective nuclear charge. Be has four protons, and N has seven protons.
Both atoms have TWO core electrons that predominately contribute to the shielding,
while the valence electrons will contribute a slight shielding effect.
So Be has an effective nuclear charge of slightly more than 2+,
and N has an effective nuclear charge of slightly  MORE THAN 5+

Thanks!